OpenStudy (anonymous):

A survey of high school juniors found that 82% of students plan on attending college. If you pick three students at random, what is the probability that at least two plan on attending college? Round to the nearest percent. I continue to get 36%, but the answer is 91%. How?

OpenStudy (anonymous):

I used this formula: \[P(x)= nC _{x}p ^{x}q ^{n-x}\] with the first 'n' as a subscript. This is the only formula included in the lesson, so I assumed it was the correct one to use on this problem.

OpenStudy (dan815):

okk i finally figured it out lol

OpenStudy (dan815):

this is how u do it!! lets look at the solution that u dont want the 3 persons u pick ___ , ____, _____ no no no = 0.18^3 no no yes = 0.18 * 0.18 * 0.82 no yes no = 0.18*0.82*0.18 yes no no = 0.82*0.18*0.18 so in other words all the ones u dont need = 0.18^3+3 * (0.18^2*0.82) now the probability of the success = just 1 - 0.18^3+3 * (0.18^2*0.82) = 91.44%

OpenStudy (dan815):

the other way to do it wud be to look at all the success so like ----, -----, ----- yes yes yes = 0.82^3 yes no yes =0.82^2*0.18 yes yes no =0.82^2*0.18 no yes yes =0.82^2*0.18 add all that up = 0.914464 or 91.44% hope this helps there is a formula u can form out of this too like P of success ^ random choices + (P of success ^ random choices - 1) * number of choices )

OpenStudy (anonymous):

Thank you so much!

OpenStudy (dan815):

np :)

OpenStudy (dan815):

also if it may help you in the future the way i arrived at the fact that those only 8 were the possible outcomes was that -----, -----,----- 2*2*2 = 8 like a coin flip except the sides have a different probability of landing. I know it doesnt sound very useful when its only for 3 people and 2 choices, but if these numbers pile up it would be a different story